The equation of a particle executing SHM is x = 3 sin (2πt+π/6), where x is in meter and t is in seconds. 

Find the amplitude, time period, maximum speed and maximum acceleration. Also find the velocity and 

acceleration at t = 0.5 s.


1
Expert's answer
2021-05-27T19:06:50-0400

a) "A=3\ m".

b)

"\omega=2\pi f=\dfrac{2\pi}{T},""T=\dfrac{2\pi}{\omega}=\dfrac{2\pi}{2\pi s^{-1}}=1\ s."

c)

"v=\dfrac{dx}{dt}=\dfrac{d}{dt}(3 sin(2\pi t+\dfrac{\pi}{6}))=3\cdot2\pi\ s^{-1} cos(2\pi t+\dfrac{\pi}{6}),""v=18.85\ \dfrac{m}{s}\cdot cos(2\pi t+\dfrac{\pi}{6})."

Maximum speed will be when "cos(2\pi t+\dfrac{\pi}{6})=1":


"v_{max}=18.85\ \dfrac{m}{s}."

d)

"a=\dfrac{dv}{dt}=\dfrac{d}{dt}(18.85\ \dfrac{m}{s}\cdot cos(2\pi t+\dfrac{\pi}{6})),""a=-18.85\ \dfrac{m}{s}\cdot2\pi\ s^{-1}\cdot sin(2\pi t+\dfrac{\pi}{6})=-118.4\ \dfrac{m}{s^2}\cdot sin(2\pi t+\dfrac{\pi}{6})."

Maximum acceleration will be when "sin(2\pi t+\dfrac{\pi}{6})=1":


"a_{max}=-118.4\ \dfrac{m}{s^2}."

e)

"v(t=0.5\ s)=18.85\ \dfrac{m}{s}\cdot cos(2\pi\ s^{-1}\cdot0.5\ s+\dfrac{\pi}{6}),""v(t=0.5\ s)=-16.32\ \dfrac{m}{s}."

f)

"a(t=0.5\ s)=-118.4\ \dfrac{m}{s^2}\cdot sin(2\pi\ s^{-1}\cdot0.5\ s+\dfrac{\pi}{6}),""a(t=0.5\ s)=59.2\ \dfrac{m}{s^2}."